OpenStudy (anonymous):

What are the coordinates of the vertex for f(x) = x2 + 6x + 13? Walk me through the steps please?

OpenStudy (michele_laino):

I can rewrite your function as below: \[\begin{gathered} y = {x^2} + 6x + 9 - 9 + 13 \hfill \\ y = {\left( {x + 3} \right)^2} + 4 \hfill \\ y - 4 = {\left( {x + 3} \right)^2} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

where did the 9's come from?

OpenStudy (anonymous):

the 4 too

OpenStudy (michele_laino):

It is a my choice in order to get a perfect square: \((x+3)^2\)

OpenStudy (anonymous):

oh i see!

OpenStudy (michele_laino):

\(4\) comes from \(-9+13=4\) In other words I added and subtracted \(9\), and such operation doesn't change the right side of your equation

OpenStudy (anonymous):

Oh okay. Thank you very much!

OpenStudy (anonymous):

next step?

OpenStudy (michele_laino):

now I maje this traslation: \[\left\{ \begin{gathered} Y = y - 4 \hfill \\ X = x + 3 \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

oops.. I made*...

OpenStudy (anonymous):

gotcha

OpenStudy (michele_laino):

sorry.. I make*

OpenStudy (michele_laino):

where \(Y,X\) are the new coordinates Using those new coordinates our equation can be rewritten as below: \[Y = {X^2}\]

OpenStudy (anonymous):

These are my answer choices: A. 2(x + 5)(x − 4) B. 2(x − 5)(x + 4) C. 2(x + 5)^2 D. 2(x + 5)(x − 5)

OpenStudy (michele_laino):

I think that those options are related to another exercise

OpenStudy (anonymous):

oh I'm sorry! you're right. gimme a sec.

OpenStudy (michele_laino):

no problem :)

OpenStudy (michele_laino):

the coordinates of the vertex of this parabola \(Y=X^2\) are: \(Y=0,X=0\)

OpenStudy (michele_laino):

substituting \(X=0,Y=0\) into the equation of my trasformation, I get: \[\left\{ \begin{gathered} y - 4 = 0 \hfill \\ x + 3 = 0 \hfill \\ \end{gathered} \right.\] please find \(x,y\)

OpenStudy (anonymous):

Here are my answer choices for real: A. (4, 4) B. (−4, 4) C. (3, 4) D. (−3, 4) and okay will do

OpenStudy (michele_laino):

for example, from the first equation, I get \(y=4\), and from the second equation, we get: \(x=...?\)

OpenStudy (anonymous):

x = -3

OpenStudy (michele_laino):

correct! so what is the right option?

OpenStudy (anonymous):

D. (-3, 4)!

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Yes! Thank you for your help!